Find the interest rate on a loan charging $765 simple interest on a principal of $3750 after 6 years. __________

Find the simple interest on the loan. $1900 at 8% for 10 years. _________

Find the term of a loan of $225 at 3.5% if the simple interest is $63. __________

Determine the amount due on the compound interest loan. (Round your answers to the nearest cent.)

$16,000 at 4% for 15 years if the interest is compounded in the following ways.

__________Annually

__________Quarterly
Answer #1

1. Simple Interest = Principal * Rate* Time/100
Rate = 765*100/(3750*6) = 3.4%

2. Simple Interest = Principal * Rate* Time/100 = 1900*8%*10/100 = 15.20

3. Simple Interest = Principal * Rate* Time/100
Time = 63*100/(3.5*225) = 8 years

4. Amount annually = 16000*(1+4%)15 = 28815.10
Amount Quarterly =  16000*(1+4%/4)15*4   =  29067.15

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well